# Fast Robot Adaptation via Hand Path Retrieval

### v3

@@ Abstract
We hand the community HAND, a simple and time-efficient method for teaching robots new manipulation tasks through human hand demonstrations. Instead of relying on task-specific robot demonstrations collected via teleoperation, HAND uses easy-to-provide hand demonstrations to retrieve relevant behaviors from task-agnostic robot play data. Using a visual tracking pipeline, HAND extracts the motion of the human hand from the hand demonstration and retrieves robot sub-trajectories in two stages: first filtering by visual similarity, then retrieving trajectories with similar behaviors to the hand. Fine-tuning a policy on the retrieved data enables real-time learning of tasks in under four minutes, without requiring calibrated cameras or detailed hand pose estimation. Experiments also show that HAND outperforms retrieval baselines by over 2×2$\times$2 × in average task success rates on real robots. Videos can be found at our project website: https://liralab.usc.edu/handretrieval/. *Equal Contribution, †Equal Advising
Keywords: Robot Adaptation, Retrieval, Play Data, Hand Demonstration
1 Introduction
Figure 1: HAND learns a policy from as little as one (1) human hand demonstration.
Robots deployed in homes, warehouses, and other dynamic, human-centric settings will need to quickly learn many tasks specified by end-users. To support this goal, robot learning algorithms for these settings must (1) scale easily across many tasks and (2) enable fast adaptation for each new task. While imitation learning has shown promise in producing capable multi-task robot policies [1, 2, 3, 4, 5], it remains difficult to scale due to its reliance on large amounts of expert-collected, task-specific teleoperation data. In contrast, task-agnostic play data—collected through free-form robot teleoperation [6, 7, 8]—is much easier to gather, as it does not require constant environment resets or task-specific labeling. The key challenge is making this data usable for teaching robots new tasks quickly. In this work, we tackle this problem by retrieving relevant robot behaviors from the play dataset using only a single human hand demonstration.
@@ 1 Introduction
We propose HAND, a simple and time-efficient approach that adapts pre-trained play policies to specific tasks using just one human hand demonstration (see Figure 1). Prior work on robot behavior retrieval relies on additional teleoperated target-task robot demonstrations [9, 10, 11, 12, 13], which are difficult for non-expert users to provide. Instead, our key insight is to extract coarse guidance from the hand demonstration—specifically, 2D relative hand motion paths—to retrieve diverse yet relevant behaviors from the play dataset.
@@
Calling back to the motivation, we aim for HAND to be scalable and fast. Towards scalability, HAND avoids the need for calibrated depth cameras [14, 15], specialized eye-in-hand setups [16], or detailed hand-pose estimation [16, 17]. Instead, it first labels a robot play dataset with 2D gripper positions relative to the RGB camera frame by tracking the gripper using a visual point-tracking model [18, 19]. When a human hand demonstration is provided, HAND tracks the hand trajectory with the same simple pipeline. The hand positions are then converted into 2D relative sub-trajectories, capturing motion independent of the starting point [20]. After an initial filtering step that removes unrelated behaviors using a visual foundation model [21], HAND retrieves matching sub-trajectories from the play dataset based on the 2D relative hand path. Finally, for fast learning, a policy pre-trained on the play dataset is LoRA-fine-tuned on the retrieved sub-trajectories, encouraging the policy to specialize in the demonstrated task. Because HAND retrieves primarily based on hand motion, it is more robust to irrelevant visual features such as background clutter and lighting changes compared to purely visual retrieval methods.
Our experiments, both in simulation in CALVIN [8] and in the real world on a WidowX robot, demonstrate that HAND enables quick adaptation to 8 diverse downstream tasks. Notably, HAND outperforms the best baseline by 2$\times$ on the real robot. We also demonstrate that HAND works with hand demos collected from completely different scenes from the robot’s. Finally, we perform a real-time learning experiment, where HAND learns a challenging long-horizon task in under 4 minutes of experiment time, from providing the hand demonstration to the trained policy, while being on average 5$\times$ faster to collect data for than robot teleoperation demos on our WidowX arm.
2 Related Works
Robot Data Retrieval.
@@ 2 Related Works
Prior work has demonstrated retrieval as an effective mechanism for extracting relevant on-robot data for training robots [9, 10, 11, 12, 13, 22]. For example, SAILOR [9] and Behavior Retrieval [10] pre-train variational auto-encoders (VAEs) on prior robot images and actions to learn a latent embedding. This latent embedding is used to retrieve states and actions from an offline dataset similar to ones provided in expert demonstration trajectories. However, retrieving based on learned full image encodings or even raw pixel values [13] can be noisy; Flow-Retrieval [11] instead trains a VAE to encode optical flows indicating movement of objects and the robot arm in the scene. Similar to Flow-Retrieval, our method HAND also retrieves based on robot arm movement. However, rather than training a dataset-specific VAE model that may not be robust to large visual differences, we retrieve from our offline robot data by primarily matching motions of a human hand demonstration using relative 2D paths of the robot end-effector in the prior data. This hand path retrieval helps us robustly retrieve relevant robot arm behaviors.
@@
STRAP [12] addresses visual retrieval robustness issues of prior work by using features from DINO-v2 [21], a large pre-trained image-input foundation model for retrieval. However, STRAP, along with all aforementioned retrieval work, assumes access to expert robot demonstrations for the target tasks. HAND  on the other hand, only requires a single, easier-to-collect human hand demonstration that results in more time-efficient learning of demonstrated tasks compared to methods requiring robot teleoperation data for retrieval. Moreover, experiments demonstrate HAND actually retrieves more task-relevant trajectories and therefore attains higher success rates compared to these methods.
Learning From Human Hands.
@@
Similar to HAND, a separate line of work has proposed methods to use human hands to learn robot policies. One approach is to train models on human video datasets to predict future object flows [23, 24] or human affordances [25, 26]. These intermediate affordance and flow representations are then used to either train a policy conditioned on this representation [23] on robot data or control a heuristic policy [24, 25, 26]. Other works focus on learning directly from human hands [14, 15, 16, 27, 17]. These works generally use hand-pose detection models aided by multiple cameras or calibrated depth cameras to convert hand poses directly to robot gripper keypoints [14, 15, 17]. However, works that exclusively retrieve human data are restricted to constrained policy representations as they must match human hand poses to robot gripper poses. Kim et al. [16] instead use an eye-in-hand camera mounted on a human demonstrator’s forearm to train an imitation learning policy conditioned on robot eye-in-hand camera observations. Unlike these prior works, HAND only requires a single RGB camera from which the robot gripper can be seen. Also, we focus on retrieving robot play data, allowing us to train arbitrarily expressive policies without constrained policy representations [14, 15, 17] or intermediate representations [23, 24, 25, 26].
3 HAND: Fast Robot Adaptation via Hand Path Retrieval
@@
We assume access to a dataset of task-agnostic robot play data, $\mathcal{D}_{\text{play}}$, consisting of trajectories ${\tau_{i}=\{(o_{t},a_{t})\}_{i=1}^{T}}$, where each $o_{t}$ is per-timestep observation that includes RGB images of the robot gripper and robot proprioceptive information, and $a_{t}$ is the robot action. These trajectories may span many scenes or tasks and can vary in length, potentially covering long-horizon behavior. We do not assume task labels (e.g., language labels), as data collection is easier to scale without labeling each sub-trajectory in a long-horizon play trajectory. [^1] HAND can also easily incorporate task labels as an extra policy conditioning input. We assume the RGB camera’s angle relative to the robot base is fixed across trajectories, which is the case for tabletop robot manipulation setups.
@@
In contrast to prior retrieval methods that rely on robot demonstrations for each target task [9, 10, 11, 12], we assume access to easy-to-provide human hand demonstrations. [^2] HAND also outperforms baselines even when they have access to robot demonstrations. For each task, a human records their hand movement without teleoperating the robot. On our real-world setup, these hand demonstrations, $\mathcal{D}_{\text{hand}}$, are on average 5×5$\times$5 × faster to collect than robot teleoperation data. Moreover, producing high-quality hand demonstrations typically requires far less effort than robot teleoperation [28, 29]. Each video in $\mathcal{D}_{\text{hand}}$ consists of a sequence of RGB images $o_{1},\ldots,o_{H}$, captured from a similar viewpoint relative to the human hand as the robot play data relative to the robot gripper.
@@
Given $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$, we aim to train a policy $\pi_{\theta}(a\mid o)$ to perform the target task demonstrated by the human in $\mathcal{D}_{\text{hand}}$. Since we do not assume task labels in $\mathcal{D}_{\text{play}}$ and we are provided no expert robot teleoperation demonstrations, we must retrieve sub-trajectories indicating how to perform the behavior demonstrated in $\mathcal{D}_{\text{hand}}$ from $\mathcal{D}_{\text{play}}$ for training $\pi$. We denote this retrieved dataset, which we later use for imitation learning, as $\mathcal{D}_{\text{retrieved}}$. Moreover, following our motivation in Section 1, we aim for our method to be fast, so that non-expert end-users can easily train the robot for many downstream tasks. Thus, the key challenges we resolve in our method HAND are: (1) designing a representation that can unify the behaviors in robot sub-trajectories and human hand demonstrations (Section 3.1), (2) retrieving relevant sub-trajectories based on a suitable distance metric between these representations (Section 3.2), and (3) time-efficiently training a policy that can perform various unseen target tasks with a high success rate without expert demonstrations (Section 3.3). See Figure 2 for an overview and Algorithm 1 for full algorithm pseudocode.
3.1 Path Distance as a Unifying Representation for Retrieval
@@
Existing robot data retrieval methods assume access to expert demonstrations from which they extract proprioceptive information (e.g., joint states and actions) alongside visual features for retrieval [9, 10, 11, 12, 13]. However, since $\mathcal{D}_{\text{hand}}$ contains only visual data and no robot actions, retrieval based purely on appearance can be noisy—especially due to the visual domain gap between hand demonstrations in $\mathcal{D}_{\text{hand}}$ and robot demonstrations in $\mathcal{D}_{\text{play}}$ (c.f., Figure 2, left). To address these issues, we propose an embodiment-agnostic, behavior-centric retrieval metric that enables matching between $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$ based on demonstrated behaviors rather than appearance.
@@
3.2 Retrieving Relevant Sub-Trajectories using Path Distance
@@
Background. For identifying relevant sub-trajectories in $\mathcal{D}_{\text{play}}$, we follow Memmel et al. [12] and use Subsequence Dynamic Time Warping (S-DTW) [31], an algorithm for aligning a shorter sequence to a portion of a longer reference sequence. Given a query sequence $Q=\{q_{1},q_{2},\dots,q_{H}\}$ and a longer reference sequence $R=\{r_{1},r_{2},\dots,r_{T}\}$, where $T>H$, the goal of S-DTW is to find a contiguous subsequence of $R$ that minimizes the total cumulative distance between elements of both sequences. In HAND, the query sequences are the 2D hand demo paths $\{(x_{t},y_{t})_{\text{hand}}\}_{t=1}^{H}$ and the reference sequences are the 2D paths generated from long-horizon robot play data $\{(x_{t},y_{t})_{\text{play}}\}_{t=1}^{T}$.
@@
Sub-Trajectory Preprocessing. To preprocess the datasets for S-DTW, we first segment the offline play dataset, $\mathcal{D}_{\text{play}}$, into variable-length sub-trajectories using a simple heuristic based on proprioception proposed in several prior works [32, 12]. In particular, we split the trajectories whenever the acceleration or velocity magnitude (depending on what proprioception data is available) drops below a predefined $\epsilon$ value, corresponding to when the teleoperator switches between tasks. We find that this simple heuristic can reasonably segment trajectories into atomic components resembling lower-level primitives. We also split the hand demonstrations evenly into smaller sub-trajectories based on how many subtasks the human operator determined they have completed. After sub-trajectory splitting, we have two sub-trajectory datasets, ${\mathcal{T}_{\text{hand}}=\{t_{1:a}^{i},t_{a:b}^{i},\dots,t_{H_{i}-|p_{\text{hand}}^{i}|:H_{i}}^{i}\forall\,\tau^{i}_{\text{hand}}\in\mathcal{D}_{\text{hand}}\}}$ and ${\mathcal{T}_{\text{play}}=\{t_{1:a}^{j},t_{a:b}^{j},\dots,t_{T_{j}-|p_{\text{play}}^{j}|:T}^{j}\forall\,\tau^{j}_{\text{play}}\in\mathcal{D}_{\text{play}}\}}$ where $|p_{\text{hand}}^{i}|$ and $|p_{\text{play}}^{j}|$ are the lengths of the last sub-trajectory paths of trajectories $i,j$ from $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$, respectively. Inspired by prior work that proposes to cluster trajectories based on relative embedding differences [20], each sub-trajectory is represented in relative 2D coordinates, i.e., ${p_{t}=[x_{t+1}-x_{t},y_{t+1}-y_{t}]}$. Relative coordinates ensure invariance based on the starting positions of the hand or gripper so that these starting positions do not influence how trajectories are retrieved.
@@
Visual Filtering. One issue with retrieving sub-trajectories based only on path distance is that different tasks can have similar movement patterns. For example, tasks like “pick up the mug” and “pick up the cube” can appear nearly identical in 2D path space. But, the retrieved trajectories for one task may not benefit learning of the other; since we don’t assume task labels in $\mathcal{D}_{\text{play}}$, a policy directly trained on “pick up the cube” retrieved sub-trajectories may still fail to pick up a mug. Therefore, before retrieving sub-trajectories with paths, we first run a visual filtering step to ensure that the sub-trajectories we retrieve will be task-relevant. We use an object-centric visual foundation model, namely DINOv2 [21], to first filter out sub-trajectories performing unrelated tasks with different objects. Specifically, we use the DINOv2 first and final frame embedding differences, representing visual object movement from the first to last frame, between human hand demos and robot play data to filter $\mathcal{T}_{\text{play}}$. We find that using this simple method is sufficient to filter out most irrelevant sub-trajectories. For a given image sequence $o_{1:H}^{\text{hand}}$ from a hand sub-trajectory and image sequence $o_{1:T}^{\text{play}}$ from a robot play sub-trajectory, we define the cost as:
$\text{C}_{\text{visual}}(o_{1:H}^{\text{hand}},o_{1:T}^{\text{play}})=\underbrace{||\text{DINO}(o_{1}^{\text{hand}})-\text{DINO}(o_{1}^{\text{play}})||_{2}^{2}}_{\text{first frame DINO embedding difference}}+\underbrace{||\text{DINO}(o_{H}^{\text{hand}})-\text{DINO}(o_{T}^{\text{play}})||_{2}^{2}}_{\text{last frame DINO embedding difference}}.$ (1)
@@
We sort these costs and take the $M$ trajectories with lowest cost as possible retrieval trajectories for each human hand demo sub-trajectory in $\mathcal{T}_{\text{hand}}$. The rest are discarded for those hand demos.
@@
3.3 Putting it All Together: Fast-Adaptation with Parameter-Efficient Policy Fine-tuning

### v4

@@ Abstract
We hand the community HAND, a simple and time-efficient method for teaching robots new manipulation tasks through human hand demonstrations. Instead of relying on task-specific robot demonstrations collected via teleoperation, HAND uses easy-to-provide hand demonstrations to retrieve relevant behaviors from task-agnostic robot play data. Using a visual tracking pipeline, HAND extracts the motion of the human hand from the hand demonstration and retrieves robot sub-trajectories in two stages: first filtering by visual similarity, then retrieving trajectories with similar behaviors to the hand. Fine-tuning a policy on the retrieved data enables real-time learning of tasks in under four minutes, without requiring calibrated cameras or detailed hand pose estimation. Experiments also show that HAND outperforms retrieval baselines by over $2\times$ in average task success rates on real robots. Videos can be found at our project website: https://liralab.usc.edu/handretrieval/.
Equal Contribution, Equal Advising
I Introduction
For robots to operate seamlessly in human-centric settings, they should be able to rapidly learn new tasks with minimal human input. Achieving this goal requires robot learning algorithms that (1) scale across many tasks and (2) adapt quickly to new ones.
Imitation learning has produced capable multi-task robot policies [1, 2, 3, 4, 5], but scaling is hindered by its reliance on vast amounts of expert-collected, task-specific teleoperation data [6]. In contrast, task-agnostic play data is far easier to collect, without requiring constant environment resets or task-specific labeling [7, 8, 9]. The challenge is in making such unstructured data usable for teaching robots new tasks quickly.
@@ I Introduction
Therefore, we propose HAND, a simple and time-efficient approach that adapts pre-trained play policies to specific tasks using just one human hand demonstration (see LABEL:fig:teaser). Unlike prior retrieval methods [10, 11, 12, 13, 14, 15] that require robot demonstrations of the target task, HAND extracts coarse, 2D relative hand motion paths from the provided human hand demonstration to guide retrieval. Thus, our approach enables even non-experts to teach robots without teleoperation.
@@
HAND enables both scalability and speed. Towards scalability, HAND avoids the need for calibrated depth cameras [16, 17], specialized eye-in-hand setups [18], or detailed hand-pose estimation [18, 19]. Instead, it first labels a robot play dataset with 2D gripper positions relative to the RGB camera frame by tracking the gripper using a visual point-tracking model [20, 21]. When a human hand demonstration is provided, HAND tracks the hand trajectory with the same simple pipeline. The hand positions are then converted into 2D relative sub-trajectories, capturing motion independent of the starting point [22]. After an initial filtering step that removes unrelated behaviors using a visual foundation model [23], HAND retrieves matching sub-trajectories from the play dataset based on the 2D relative hand path. Finally, towards speed, a policy pre-trained on the play dataset is LoRA-fine-tuned on the retrieved sub-trajectories, encouraging the policy to specialize in the demonstrated task. Because HAND retrieves primarily based on hand motion, it is robust to irrelevant visual features such as background clutter and lighting changes compared to purely visual retrieval methods.
[…]
II Related Works
@@
Robot Data Retrieval. Prior work has demonstrated retrieval as an effective mechanism for extracting relevant on-robot data for training robots [10, 11, 12, 13, 14, 24, 15, 25]. For example, SAILOR [10] and Behavior Retrieval [11] pre-train variational auto-encoders (VAEs) on prior robot images and actions to learn a latent embedding. This latent embedding is used to retrieve states and actions from an offline dataset similar to ones provided in expert demonstration trajectories. However, retrieving based on learned full image encodings or even raw pixel values [14] can be noisy; Flow-Retrieval [12] instead trains a VAE to encode optical flows indicating movement of objects and the robot arm in the scene. Similar to Flow-Retrieval, our method HAND also retrieves based on robot arm movement. However, rather than training a dataset-specific VAE model that may not be robust to large visual differences, we retrieve from our offline robot data by primarily matching motions of a human hand demonstration using relative 2D paths of the robot end-effector in the prior data. This hand path retrieval helps us robustly retrieve relevant robot arm behaviors.
@@
STRAP [13] addresses visual retrieval robustness issues of prior work by using features from DINO-v2 [23], a large pre-trained image-input foundation model for retrieval. However, STRAP, along with all aforementioned retrieval work, assumes access to expert robot demonstrations for the target tasks. HAND, on the other hand, only requires a single, easier-to-collect human hand demonstration that results in more time-efficient learning of demonstrated tasks compared to methods requiring robot teleoperation data for retrieval. Moreover, experiments demonstrate HAND actually retrieves more task-relevant trajectories and therefore attains higher success rates compared to these methods.
@@
Learning From Human Hands. Similar to HAND, a separate line of work has proposed methods to use human hands to learn robot policies. One approach is to train models on human video datasets to predict future object flows [26, 27] or human affordances [28, 29]. These intermediate affordance and flow representations are then used to either train a policy conditioned on this representation [26] on robot data or control a heuristic policy [27, 28, 29]. Other works focus on learning directly from human hands [16, 17, 18, 30, 19]. These works generally use hand-pose detection models aided by multiple cameras or calibrated depth cameras to convert hand poses directly to robot gripper keypoints [16, 17, 19]. However, works that exclusively retrieve human data are restricted to constrained policy representations as they must match human hand poses to robot gripper poses. [18] instead use an eye-in-hand camera mounted on a human demonstrator’s forearm to train an imitation learning policy conditioned on robot eye-in-hand camera observations. Unlike these prior works, HAND only requires a single RGB camera from which the robot gripper can be seen. Also, we focus on retrieving robot play data, allowing us to train arbitrarily expressive policies without constrained policy representations [16, 17, 19] or intermediate representations [26, 27, 28, 29].
III HAND: Fast Robot Adaptation
via Hand Path Retrieval
@@ III HAND: Fast Robot Adaptation
We assume access to a dataset of task-agnostic robot play data, $\mathcal{D}_{\text{play}}$, consisting of trajectories $\tau_{i}=\{(o_{t},a_{t})\}_{i=1}^{T}$}, where each $o_{t}$ is per-timestep observation that includes RGB images of the robot gripper and robot proprioceptive information, and $a_{t}$ is the robot action. These trajectories may span many scenes, tasks, and time horizons. We do not assume task labels (e.g., language labels), as data collection is easier to scale without labeling each sub-trajectory in a long-horizon play trajectory. [^1] Section IV demonstrates that HAND can also incorporate language labels as extra policy conditioning.
@@
In contrast to retrieval methods that rely on robot demonstrations for each target task [10, 11, 12, 13], we assume access to easy-to-provide human hand demonstrations. For each task, a human records their hand movement without teleoperating the robot. On our real-world setup, these hand demonstrations, $\mathcal{D}_{\text{hand}}$, are on average $5\times$ faster to collect than robot teleoperation data. Moreover, producing high-quality hand demonstrations typically requires far less effort than robot teleoperation [31, 32]. Each video in $\mathcal{D}_{\text{hand}}$ consists of a sequence of RGB images $o_{1},\ldots,o_{H}$, captured from a similar viewpoint relative to the human hand as trajectories in the robot play data relative to the robot gripper. [^2] Section IV demonstrates HAND works under large camera angle shifts.
@@
Given $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$, we aim to train a policy $\pi_{\theta}(a\mid o)$ to perform the target task demonstrated by the human in $\mathcal{D}_{\text{hand}}$. Since we do not assume task labels in $\mathcal{D}_{\text{play}}$ and we are provided no expert robot teleoperation demonstrations, we must retrieve sub-trajectories indicating how to perform the behavior demonstrated in $\mathcal{D}_{\text{hand}}$ from $\mathcal{D}_{\text{play}}$ for training $\pi$. We denote this retrieved dataset, later used for imitation learning, as $\mathcal{D}_{\text{retrieved}}$. Moreover, following our motivation in Section I, we aim for our method to be fast, so that non-expert end-users can easily train the robot for many downstream tasks.
Thus, the key challenges we resolve in our method HAND are: (1) designing a representation that can unify the behaviors in robot sub-trajectories and human hand demonstrations (Section III-B), (2) retrieving relevant sub-trajectories based on a suitable distance metric between these representations (Section III-C), and (3) quickly training a policy that can perform various unseen target tasks with a high success rate without expert demonstrations (Section III-D). See Figure 2 for an overview.
III-B Path Distance as a Unifying Representation for Retrieval
@@
Prior robot retrieval methods assume access to expert demonstrations from which they extract proprioceptive information (e.g., joint angles and actions) alongside visual features for retrieval [10, 11, 12, 13, 14]. However, since $\mathcal{D}_{\text{hand}}$ contains only visual data and no robot actions, retrieval based purely on appearance can be noisy—especially due to the visual domain gap between hand demonstrations in $\mathcal{D}_{\text{hand}}$ and robot demonstrations in $\mathcal{D}_{\text{play}}$ (c.f., Figure 2, left). To address these issues, we propose an embodiment-agnostic, behavior-centric retrieval metric that enables matching between $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$ based on demonstrated behaviors rather than appearance.
@@
III-C Retrieving Relevant Sub-Trajectories using Path Distance
@@
Background. For identifying relevant sub-trajectories in $\mathcal{D}_{\text{play}}$, we use Subsequence Dynamic Time Warping (S-DTW) [34], an algorithm for aligning a shorter sequence to a portion of a longer reference sequence prior work has demonstrated effective for sub-trajectory retrieval [13]. Given a query sequence $Q=\{q_{1},q_{2},\dots,q_{H}\}$ and a longer reference sequence $R=\{r_{1},r_{2},\dots,r_{T}\}$, where $T>H$, the goal of S-DTW is to find a contiguous subsequence of $R$ that minimizes the total cumulative distance between elements of both sequences. In HAND, the query sequences are the 2D hand demonstration paths $\{(x_{t},y_{t})_{\text{hand}}\}_{t=1}^{H}$ and the reference sequences are the 2D paths generated from long-horizon robot play data $\{(x_{t},y_{t})_{\text{play}}\}_{t=1}^{T}$.
@@
Sub-Trajectory Preprocessing. To preprocess the datasets for S-DTW, we first segment the offline play dataset, $\mathcal{D}_{\text{play}}$, into variable-length sub-trajectories using a simple heuristic based on proprioception proposed in several prior works [35, 13]. In particular, we split the trajectories whenever the acceleration or velocity magnitude (depending on what proprioception data is available) drops below a predefined $\epsilon$ value, corresponding to when the teleoperator switches between tasks. We find that this simple heuristic can reasonably segment trajectories into atomic components resembling lower-level primitives. We also split the hand demonstrations evenly into smaller sub-trajectories based on how many subtasks the human operator determined they have completed. After sub-trajectory splitting, we have two sub-trajectory datasets, $\mathcal{T}_{\text{hand}}=\{t_{1:a}^{i},t_{a:b}^{i},\dots,t_{H_{i}-|p_{\text{hand}}^{i}|:H_{i}}^{i}\forall\,\tau^{i}_{\text{hand}}\in\mathcal{D}_{\text{hand}}\}$} and $\mathcal{T}_{\text{play}}=\{t_{1:a}^{j},t_{a:b}^{j},\dots,t_{T_{j}-|p_{\text{play}}^{j}|:T}^{j}\forall\,\tau^{j}_{\text{play}}\in\mathcal{D}_{\text{play}}\}$} where $|p_{\text{hand}}^{i}|$ and $|p_{\text{play}}^{j}|$ are the lengths of the last sub-trajectory paths of trajectories $i,j$ from $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$, respectively. Finally, each sub-trajectory is represented in relative 2D coordinates, i.e., $p_{t}=[x_{t+1}-x_{t},y_{t+1}-y_{t}]$}. Relative coordinates ensure retrieval invariance to the initial positions of the hand or gripper [22].
@@
Visual Filtering. One issue with retrieving sub-trajectories based only on path distance is that different tasks can have similar movement patterns. For example, tasks like “pick up the mug” and “pick up the cube” can appear nearly identical in 2D path space [4]. But, the retrieved trajectories for one task may not benefit learning of the other; since we do not assume task labels in $\mathcal{D}_{\text{play}}$, a policy directly trained on “pick up the cube” retrieved sub-trajectories may still fail to pick up a mug. Therefore, before retrieving sub-trajectories with paths, we first run a visual filtering step to ensure that the sub-trajectories we retrieve will be task-relevant. We use an object-centric visual foundation model, namely DINOv2 [23], to first filter out sub-trajectories performing unrelated tasks with different objects. Specifically, we use the DINOv2 first and final frame embedding differences, representing visual object movement from the first to last frame, between human hand demonstrations and robot play data to filter $\mathcal{T}_{\text{play}}$. We find that using this simple method is sufficient to filter out most irrelevant sub-trajectories. For a given image sequence $o_{1:H}^{\text{hand}}$ from a hand sub-trajectory and image sequence $o_{1:T}^{\text{play}}$ from a robot play sub-trajectory, we define the cost as:
[…]
@@
We take the $M$ trajectories with lowest cost as possible retrieval trajectories from $\mathcal{D}_{\text{play}}$ for each human demonstration sub-trajectory in $\mathcal{T}_{\text{hand}}$. The rest are ignored for those hand demonstrations.
@@
III-D Putting it All Together: Fast-Adaptation with Parameter-Efficient Policy Fine-tuning